OpenStudy (anonymous):

A manufacturer of small glass figurines discovers that it costs $ 1900 for a production run of 1000 and $ 2100 for a production run of 1300 . Assuming that cost is a linear function of the number of items, find the overhead and the marginal cost of a figurine. The overhead is $ . the marginal cost of a figurine is $ .

OpenStudy (slaaibak):

Marginal cost = (2100 - 1900)/(1300 - 1000) = 200/300 = 2/3 So Overhead costs = 1900 - 1000 * 2/3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thansk